HYPERFINE STRUCTURE IN THE PURE ROTATIONAL SPECTRUM OF ^{208}Pb^{35}Cl

CHRISTOPHER T. DEWBERRY, GARRY S. GRUBBS II, KERRY C. ETCHISON AND \underline{STEPHEN A. COOKE}, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF NORTH TEXAS, 1155 UNION CIRCLE #305070, DENTON, TX 76203-5017, U.S.A.

Initially in our laboratory the pure rotational spectrum of the title molecule was studied using a Balle-Flygare Fourier transform microwave spectrometer. Analysis was troublesome and so the spectrum was remeasured using a chirped pulse Fourier transform microwave (CP-FTMW) spectrometer. The correct intensity aspect of the CP-FTMW experiment allowed succesful quantum number assignments for the hyperfine structure for the correct isotopologue. Spectroscopic constants have been obtained from a fit to a data set consisting of our measurements combined with those of a prior study on the X_{2}^{2}Π_{\frac{3}{2}} → X_{1}^{2}Π_{\frac{1}{2}} fine structure transitions^{a}.

aK. Ziebarth, K. D. Setzer, O. Shestakov and E. H. Fink J. Mol. Spectrosc. \underline{191} 108, 1998.